In my class today, we learned rates. Here is one of my problems: 55mi/h=?ft/s I have no clue what to do. I wrote out 55mi * ____ = ____ft
1 h s
What am I supposed to do next?

Answer by rchill(405)

There are several ways to solve this. Let's look at it logically. The speed of 55mph means that in an hour we travel 55*5280=290400 feet since there are 5280 feet in one mile. There are 60*60=3600 seconds in one hour, so to compute how far we travel in one of those seconds: 290400/3600=80.67 feet.